Morrisey & Brown, Ltd., of Sydney is a merchandising company that is the sole distributor of a product that is increasing in popularity among Australian consumers. The company’s income statements for the three most recent months follow:

Morrisey & Brown, Ltd.
Income Statements
For the Three Months Ended September 30
July August September
Sales in units 5,250 5,750 6,250
Sales $ 630,000 $ 690,000 $ 750,000
Cost of goods sold 378,000 414,000 450,000
Gross margin 252,000 276,000 300,000
Selling and administrative expenses:
Advertising expense 30,600 30,600 30,600
Shipping expense 46,500 49,500 52,500
Salaries and commissions 93,000 99,500 106,000
Insurance expense 10,200 10,200 10,200
Depreciation expense 21,300 21,300 21,300
Total selling and administrative expenses 201,600 211,100 220,600
Net operating income $ 50,400 $ 64,900 $ 79,400

Required:

1. By analyzing the data from the company's income statements, classify each of its expenses (including cost of goods sold) as either variable, fixed, or mixed.

2. Using the high-low method, separate each mixed expense into variable and fixed elements. Express the variable and fixed portions of each mixed expense in the form Y = a + bX.

3. Redo the company’s income statement at the 6,250-unit level of activity using the contribution format.
Answer #1

1.

Cost Type Cost Behavior
Cost of Goods Sold Variable Cost at $72 per unit
Advertising Expense Fixed Cost
Shipping Expense Mixed Cost
Salaries and Commissions Mixed Cost
Insurance Expense Fixed
Depreciation Expense Fixed

2) Shipping Expense:

Variable Cost = (52,500 - 46,500) / (6250 - 5250) = $6per unit

Fixed Cost = 52500 - (6250 x 6) = $15,000

Cost formula for Shipping Expense: Y = 15,000 + 6X

Salaries and Commissions:

Variable Cost = (106,000 - 93,000) / (6250 - 5250) = $13per unit

Fixed Cost = 106,000 - (6250 x 13) = $24,750

Cost formula for Salaries and Commissions: Y = 24,750 + 13X

3.

Sales Revenue 7,50,000
less Variable Costs
Cost of goods sold 4,50,000 (6250 x 72)
Variable Shipping Expense 37,500(6250 x 6)
Variable Salaries and Commissions 81,250 (6250 x 13)
Contribution Margin 1,81,250
less Fixed Costs
Advertising Expense 30,600
Fixed Shipping Expense 15,000
Fixed Salaries and Commissions 24,750
Insurance Expense 10,200
Depreciation Expense 21,300
Net Operating Income 79400